OpenStudy (anonymous):

How many ways are there to choose a committee of 3 people from a group of 9 people?

jimthompson5910 (jim_thompson5910):

Hint: use the formula n C r = (n!)/(r!*(n-r)!) In this case, n = 9 r = 3

OpenStudy (anonymous):

i have no clue what the ~!~ mean??

jimthompson5910 (jim_thompson5910):

example 6! = 6*5*4*3*2*1

jimthompson5910 (jim_thompson5910):

it's factorial notation and hopefully you've learned about it before

OpenStudy (anonymous):

i still dont understand... i have never learned about this

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 9 C 3 = (9!)/(3!*(9-3)!) 9 C 3 = (9!)/(3!*6!) 9 C 3 = (9*8*7*6!)/(3!*6!) 9 C 3 = (9*8*7)/(3!) 9 C 3 = (9*8*7)/(3*2*1) 9 C 3 = (504)/(6) 9 C 3 = 84 So there are 84 ways to form a committee If this all looks foreign, then you must have missed part of the lesson.

jimthompson5910 (jim_thompson5910):

My suggestion is that if you've never heard of factorials, then it would be best to start here http://www.purplemath.com/modules/factorial.htm and that hopefully will make more sense

OpenStudy (anonymous):

thankyou!

jimthompson5910 (jim_thompson5910):

you're welcome